# SpGesture: Source-Free Domain-adaptive sEMG-based Gesture Recognition with Jaccard Attentive Spiking Neural Network

### v2

@@ Abstract
Surface electromyography (sEMG) based gesture recognition offers a natural and intuitive interaction modality for wearable devices. Despite significant advancements in sEMG-based gesture-recognition models, existing methods often suffer from high computational latency and increased energy consumption. Additionally, the inherent instability of sEMG signals, combined with their sensitivity to distribution shifts in real-world settings, compromises model robustness. To tackle these challenges, we propose a novel SpGesture framework based on Spiking Neural Networks, which possesses several unique merits compared with existing methods: (1) Robustness: By utilizing membrane potential as a memory list, we pioneer the introduction of Source-Free Domain Adaptation into SNN for the first time. This enables SpGesture to mitigate the accuracy degradation caused by distribution shifts. (2) High Accuracy: With a novel Spiking Jaccard Attention, SpGesture enhances the SNNs’ ability to represent sEMG features, leading to a notable rise in system accuracy. To validate SpGesture’s performance, we collected a new sEMG gesture dataset which has different forearm postures, where SpGesture achieved the highest accuracy among the baselines ($89.26\%$). Moreover, the actual deployment on the CPU demonstrated a system latency below 100ms, well within real-time requirements. This impressive performance showcases SpGesture’s potential to enhance the applicability of sEMG in real-world scenarios. The code is available at https://github.com/guoweiyu/SpGesture/.
@@ 1 Introduction
In recent years, Spiking Neural Networks (SNNs) [41, 21, 20] provide an unparalleled chance for developing more practical and efficient sEMG-based gesture recognition systems. SNNs emulate the spiking behavior of biological neurons with a unique binary information communication protocol [53]. This binary communication is particularly amenable to the architectural specifics of sparse neuromorphic hardware [50]. Besides, the primary computations in SNNs revolve around spike-based accumulate (AC) operations [13]. The event-driven nature of these networks [51, 23, 82] enables calculations to be made only when there is a change or ‘event’ in the input, thereby circumventing the need to process zero values. Therefore, compared to conventional Artificial Neural Networks (ANNs) that typically rely on energy-demanding multiply-and-accumulate (MAC) operations [64] and require traditional dense computing hardware like GPUs, SNNs demonstrate substantially lower power consumption [42], positioning them as a promising candidate for developing energy-efficient gesture recognition systems [2, 71, 52].
Although SNNs are computationally efficient, they struggle to match the accuracy of ANN-based models [15]. In particular, the major problem is that the binary and sparse feature representations make it difficult to perform regular contiguous similarity computations. This limitation hinders expressive operations like attention mechanisms and advanced representation alignment algorithms, such as domain adaptation. For example, attention-based structures like the Transformer have demonstrated remarkable performance in Natural Language Processing [16, 40, 8, 69], Computer Vision [26, 30, 39, 70, 68], Time-Series Processing [76, 77, 75] and Decision Making tasks[56, 57], leading to a wave of attention-centric architecture designs, underscoring the importance and versatility of attention mechanisms in deep learning. However, with the proportion of ‘1’s typically less than $5\%$, the dot product in cosine similarity inherent to attention mechanisms tends to yield results close to zero [67]. Existing work often first converts spike signals into continuous values for similarity calculations, but this can increase the inference latency and energy consumption of SNNs. There is still a lack of methods for directly implementing advanced operations on spike features in SNN for realizing effective sEMG-based gesture recognition systems.
@@
  • We propose a Jaccard-based attention mechanism specifically designed for SNNs. This innovative approach preserves the original computational characteristics of SNNs, boosts inference efficiency, and counteracts the accuracy degradation caused by sparse spiking sequences.
@@ 3 Preliminaries
Data Collection: In human-computer interaction studies involving surface electromyography (sEMG), diverse and representative data sets are crucial. Traditional research often collects sEMG data from a single forearm posture [4, 31], but variations in forearm posture significantly influence sEMG data distribution. Our methodology incorporates gestures performed in different forearm postures to better reflect real-world conditions. Participants were instructed to replicate gestures and forearm postures shown on a screen. Our dataset includes ten gestures across three forearm postures: P1 (forearm horizontal on a surface), P2 (forearm elevated diagonally with elbow anchored), and P3 (forearm horizontal). Each gesture was held for five seconds with a five-second relaxation period, repeated six times per posture. This approach aims to provide robust sEMG data reflecting practical variability. For further details on dataset collection, including information about the acquisition devices and specific measures taken, please refer to appendix A.1.
@@ 4.1.1 Network Overview
Our proposed Jaccard Attentive Spiking Neural Network (JASNN) comprises four primary components: a Convolutional Leaky Integrate-and-Fire (ConvLIF) based spike encoder and feature extractor, a Spiking Jaccard attention mechanism, LIF-based Classifier, and a membrane potential recording module. We detailed the first the third components in appendix A.4 A.6.
@@ 4.1.2 Spiking Jaccard Attention
Attention mechanisms have enhanced DNNs in time-series analysis by focusing on important temporal aspects for better predictions. However, applying attention mechanisms to Spiking Neural Networks (SNNs) presents unique challenges. Firstly, SNNs’ sparse neuron activation (about 2%) makes the dot product operation in attention mechanisms produce sparse spike trains, hindering learning due to reduced signal strength. Secondly, using the softmax function for attention scores increases computational complexity and energy consumption, which is unsuitable for SNNs’ low-power, efficient processing requirements.
@@
Figure 2: Comparison of MA-SNN and Spiking Jaccard Attention Modules MA-SNN [67] uses fully connected layers with pooling but lacks a querying mechanism, leading to continuous intermediate values and lower efficiency. Our Spiking Jaccard Attention uses spike values for intermediate representations, enhancing efficiency and accuracy.
@@
This approach enables efficient computation of the Jaccard similarity by taking advantage of the sparsity of the data in SNNs. By computing the sums over the element-wise minimum and maximum operations instead of using matrix dot multiplication operations, our algorithm becomes more easily deployable on spiking chips, thereby enhancing the computational efficiency of the attention mechanism within SNNs. We add a tiny constant to the denominator to avoid a division by zero when there are no spikes in the spike train.
@@ 4.2 Spiking Source-Free Domain Adaptation based on Membrane Potential Memory
Another challenge is that the intermediate layer features in SNNs are represented by Spike Trains, and existing methods for finding neighbors cannot directly compute them. To identify the semantically closest neighbors to a target domain sample, we utilize the membrane potential from the Memory Layer to construct a Membrane Potential Memory List. Note that we only use target source data to generate Membrane Potential Memory List. The membrane potential encapsulates both spatial and temporal features, rendering it a more informative and efficient tool for our purpose. Membrane Potential Memory $M_{n}=\left\{V^{m}_{n,t}\right\}_{t=1}^{T}$ can be computed by:
@@ 5.1 Gesture Recognition based on sEMG
In terms of Top-1 Accuracy, our JASNN model, which integrates the SNN framework with the SJA mechanism, outperforms other DNN models, including CNN, TCN [5], Transformer [60], GRU [11], Informer [80], and a hybrid TCN with an Attention mechanism. This superior performance is due to: (1) The SNN structure’s alignment with the biological basis of sEMG generation, providing a natural modeling of the processes. (2) The SJA mechanism’s enhancement of sparse spike train features, focusing on the key characteristics of sEMG signals. Compared to other SNN models like LSNN [7], SIB+SNN [78], and SCNN, our model achieves higher accuracy. Models like SIB+SNN and SCNN perform lower, likely due to the absence of a feature enhancement design like SJA, which is crucial for capturing the temporal dynamics of sEMG signals. Incorporating SJA into Xu et al.‘s LSNN network [63] significantly improved performance, demonstrating SJA’s scalability in recurrent SNNs.
@@ 5.4 Result of Spiking Source-Free Domain Adaptation
In our investigation, we used the same dataset to experiment with three different methodologies, namely Pseudo-Label [34, 17] method, Neighborhood-guided Pseudo-Labels [28] method, and our proposed Probabilistic Label Generation method. In this experiment, the Pseudo-Label method determines the pseudo-label by taking the mode of the $k$ samples. The Neighborhood-guided Pseudo-Labels method, conversely, involves choosing the nearest $k$ samples from the Memory List, then randomly selecting one from these $k$ samples as the pseudo-label. Details of our proposed method have been elaborated on in the previous sections of the paper.
@@
Figure 4(a) and 4(b) represent the performance variations when deploying the model trained on Posture 1 to Posture 2 and 3, respectively, both with and without the use of our SSFDA. This is portrayed using a violin plot. It can be observed that the use of SSFDA indeed shifts the distribution of accuracy across different subjects upward as a whole. Particularly, our method exhibits superior performance after applying SSFDA compared to the other two methods. Furthermore, the standard deviation of performance across various subjects is minimal for our method, demonstrating the robustness of our methodology when employing SSFDA. We detailed the differences by individuals in 10(a) 10(b).
@@ A.1 Details of Data Collection
The experiments were carried out using the DataLITE wireless LE230 and DataLITE PIONEER, commercial sEMG acquisition systems from Biometrics Ltd. [^1] https://www.biometricsltd.com/ The device’s sampling rate is 2000Hz, allowing for high-resolution data capture of the electrical activities in the muscles during the performance of gestures. Eight LE230 sEMG sensors were uniformly and equidistantly affixed to the surface of the participant’s right forearm.
@@ A.2 Details of Data Preprocessing
$\mathrm{RMS}(\mathbf{X})=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}},$ (11)
@@ A.4 Conv-based Leaky Integrate-and-Fire Neurons
To transform the input into a spike train and learn the spike representation of sEMG samples, we use a smaller quantity of Conv-based Leaky Integrate-and-Fire (LIF) neurons. As shown in Figure 8(a)(a), Conv-based LIF neurons can process spatial and temporal features concurrently. Since sEMG signals consist of pulses from multiple motor neurons at the skin’s surface combined with noise, using LIF neurons is biologically plausible for decoding these signals into spikes. Conv-LIF neurons also constrain the decoding range within local time, mitigating the impact of noise on global decoding. This method aligns with the biological characteristics of sEMG signals and effectively manages noise interference.
Initially, convolution is used to extract spatial features $S^{t,n}$ from the present spatial input $X^{t}$. These extracted features are then combined with the temporal features $T^{t-1,n}$ of the previous moment to form the current membrane potential $M^{t}$. Following this, the Fire and Leak module, based on the current membrane potential, generates the temporal features $T^{t,n}$ for the subsequent moment and the spatial features $S^{t,n+1}$for the next layer of the network. A significant event, known as a spike, occurs if the membrane potential $M^{t}$ exceeds a threshold value, denoted as $V_{th}$. In such an event, $M^{t}$ is reset to a value $V_{reset}$, and $T^{t,n}$, the post-synaptic potential, is calculated as the product of $S^{t,n+1}$ and $V_{reset}$. On the contrary, if the membrane potential $M^{t}$ is lesser than the threshold $V_{th}$, a spike does not occur. In this case, $T^{t,n}$ is determined by $M^{t}$ and is computed as the product of the time constant $e^{-dt/\tau}$ and $M^{t}$.
@@ A.6 Details of LIF-based Classifier
$\begin{cases}M^{t}=T^{t-1}+X^{t},\\ S^{t}=Hea\left(M^{t}-u_{th}\right),\\ T^{t}=V_{reset}S^{t}+\left(e^{-\frac{dt}{\tau}}M^{t}\right)\odot\left(1-S^{t}\right).\end{cases}$ (13)
In these equations, $M^{t}$ represents the membrane potential at the $t$-th time step. $X^{t}$ represents the spatial feature and $T^{t-1}$ represents the temporal feature. $\odot$ stands for element-wise multiplication. When $M^{t}$ reaches or exceeds the threshold $u_{th}$, it is reset to $V_{reset}$, and a spike ($S^{t}$ = 1) is emitted. Otherwise, $M^{t}$ evolves according to the given dynamics, and no spike is emitted ($S[n]=0$). $e^{-dt/\tau}<1$ stands for the decay factor.

### v3

@@ Abstract
Surface electromyography (sEMG) based gesture recognition offers a natural and intuitive interaction modality for wearable devices. Despite significant advancements in sEMG-based gesture recognition models, existing methods often suffer from high computational latency and increased energy consumption. Additionally, the inherent instability of sEMG signals, combined with their sensitivity to distribution shifts in real-world settings, compromises model robustness. To tackle these challenges, we propose a novel SpGesture framework based on Spiking Neural Networks, which possesses several unique merits compared with existing methods: (1) Robustness: By utilizing membrane potential as a memory list, we pioneer the introduction of Source-Free Domain Adaptation into SNN for the first time. This enables SpGesture to mitigate the accuracy degradation caused by distribution shifts. (2) High Accuracy: With a novel Spiking Jaccard Attention, SpGesture enhances the SNNs’ ability to represent sEMG features, leading to a notable rise in system accuracy. To validate SpGesture’s performance, we collected a new sEMG gesture dataset which has different forearm postures, where SpGesture achieved the highest accuracy among the baselines ($89.26\%$). Moreover, the actual deployment on the CPU demonstrated a latency below 100ms, well within real-time requirements. This impressive performance showcases SpGesture’s potential to enhance the applicability of sEMG in real-world scenarios. The code is available at https://github.com/guoweiyu/SpGesture/.
@@ 1 Introduction
In recent years, Spiking Neural Networks (SNNs) [41, 21, 20] provide an unparalleled chance for developing more practical and efficient sEMG-based gesture recognition systems. SNNs emulate the spiking behavior of biological neurons with a unique binary information communication protocol [53]. This binary communication is particularly amenable to the architectural specifics of sparse neuromorphic hardware [50]. Besides, the primary computations in SNNs revolve around spike-based accumulate (AC) operations [13]. The event-driven nature of these networks [51, 23, 82] enables calculations to be made only when there is a change or ‘event’ in the input, thereby circumventing the need to process zero values. Therefore, compared to conventional Artificial Neural Networks (ANNs) that typically rely on energy-demanding multiply-and-accumulate (MAC) operations [64] and are normally deployed on high-computing-power hardware like GPUs, SNNs demonstrate substantially lower power consumption [42], positioning them as a promising candidate for developing energy-efficient gesture recognition systems [2, 71, 52].
Although SNNs are computationally efficient, they struggle to match the accuracy of ANN-based models [15]. In particular, the major problem is that the binary and sparse feature representations make it difficult to perform regular contiguous similarity computations. This limitation hinders expressive operations like attention mechanisms and advanced representation alignment algorithms, such as domain adaptation. For example, attention-based structures like the Transformer have demonstrated remarkable performance in Natural Language Processing [16, 40, 8, 69], Computer Vision [26, 30, 39, 70, 68], Time-Series Processing [76, 77, 75] and Decision-Making tasks[56, 57], leading to a wave of attention-centric architecture designs, underscoring the importance and versatility of attention mechanisms in deep learning. However, with the proportion of ‘1’s typically less than $5\%$, the dot product in cosine similarity inherent to attention mechanisms tends to yield results close to zero [67]. Existing work often first converts spike signals into continuous values for similarity calculations, but this can increase the inference latency and energy consumption of SNNs. There is still a lack of methods for directly implementing advanced operations on spike features in SNN for realizing effective sEMG-based gesture recognition systems.
@@
  • We propose a Jaccard similarity-based attention mechanism specifically designed for SNNs. This innovative approach preserves the original computational characteristics of SNNs, boosts inference efficiency, and counteracts the accuracy degradation caused by sparse spiking sequences.
@@ 3 Preliminaries
Data Collection: In human-computer interaction studies involving sEMG, diverse and representative data sets are crucial. Traditional research often collects sEMG data from a single forearm posture [4, 31], but variations in forearm posture significantly influence sEMG data distribution. Our methodology incorporates gestures performed in different forearm postures to better reflect real-world conditions. Participants were instructed to replicate gestures and forearm postures shown on a screen. Our dataset includes ten gestures across three forearm postures: P1 (forearm horizontal on a surface), P2 (forearm elevated diagonally with elbow anchored), and P3 (forearm horizontal). Each gesture was held for five seconds with a five-second relaxation period, repeated six times per posture. This approach aims to provide robust sEMG data reflecting practical variability. For further details on dataset collection, including information about the acquisition devices and specific measures taken, please refer to appendix A.1.
@@ 4.1.1 Network Overview
Our proposed Jaccard Attentive Spiking Neural Network (JASNN) comprises four primary components: a Convolutional Leaky Integrate-and-Fire (ConvLIF) based spike encoder and feature extractor, a Spiking Jaccard attention mechanism, LIF-based Classifier, and a membrane potential recording module. We detailed the first and third components in appendix A.4 A.6.
@@ 4.1.2 Spiking Jaccard Attention
Attention mechanisms have enhanced DNNs in time-series analysis by focusing on important temporal aspects for better predictions. However, applying attention mechanisms to Spiking Neural Networks (SNNs) presents unique challenges. Firstly, SNNs’ sparse neuron activation makes the dot product operation in attention mechanisms produce sparse spike trains, hindering learning due to reduced signal strength. Secondly, using the softmax function for attention scores increases computational complexity and energy consumption, which is unsuitable for SNNs’ efficient processing requirements.
@@
Figure 2: Comparison of MA-SNN and Spiking Jaccard Attention Modules. MA-SNN [67] uses fully connected layers with pooling but lacks a querying mechanism, leading to continuous intermediate values and lower efficiency. Our Spiking Jaccard Attention uses spike values for intermediate representations, enhancing efficiency and accuracy.
@@
This approach enables efficient computation of the Jaccard similarity by taking advantage of the sparsity of the data in SNNs. By computing the sums over the element-wise minimum and maximum operations instead of using matrix dot multiplication operations, our algorithm becomes more easily deployable on Neuromorphic chips, thereby enhancing the computational efficiency of the attention mechanism within SNNs. We add a tiny constant to the denominator to avoid a division by zero when there are no spikes in the spike train.
@@ 4.2 Spiking Source-Free Domain Adaptation based on Membrane Potential Memory
Another challenge is that the intermediate layer features in SNNs are represented by Spike Trains, and existing methods for finding neighbors cannot directly compute them. To identify the semantically closest neighbors to a target domain sample, we utilize the membrane potential from the Memory Layer to construct a Membrane Potential Memory List. Note that we only use target source data to generate the Membrane Potential Memory List. The membrane potential encapsulates both spatial and temporal features, rendering it a more informative and efficient tool for our purpose. Membrane Potential Memory $M_{n}=\left\{V^{m}_{n,t}\right\}_{t=1}^{T}$ can be computed by:
@@ 5.1 Gesture Recognition based on sEMG
In terms of Top-1 Accuracy, our JASNN model, which integrates the SNN framework with the SJA mechanism, outperforms other DNN models, including CNN, TCN [5], Transformer [60], GRU [11], Informer [80], and a hybrid TCN with an Attention mechanism. This superior performance is due to: 1) The SNN structure’s alignment with the biological basis of sEMG generation, providing a natural modeling of the processes. 2) The SJA mechanism’s enhancement of sparse spike train features, focusing on the key characteristics of sEMG signals. Compared to other SNN models like LSNN [7], SIB+SNN [78], and SCNN, our model achieves higher accuracy. Models like SIB+SNN and SCNN perform lower, likely due to the absence of a feature enhancement design like SJA, which is crucial for capturing the temporal dynamics of sEMG signals. Incorporating SJA into Xu et al.‘s LSNN network [63] significantly improved performance, demonstrating SJA’s scalability in recurrent SNNs.
@@ 5.4 Result of Spiking Source-Free Domain Adaptation
In our investigation, we used the same dataset to experiment with three different methodologies, namely Pseudo-Label [34, 17] method, Neighborhood-guided Pseudo-Labels [28] method, and our proposed Probabilistic Label Generation method. In this experiment, the Pseudo-Label method determines the pseudo-label by taking the mode of the $k$ samples. Conversely, the neighborhood-guided Pseudo-labeles method involves choosing the nearest $k$ samples from the memory list and then randomly selecting one from these $k$ samples as the pseudo-label. Details of our proposed method have been elaborated on in the previous sections of the paper.
@@
Figure 4(a) and 4(b) represent the performance variations when deploying the model trained on Posture 1 to Posture 2 and 3, respectively, both with and without the use of our SSFDA. This is portrayed using a violin plot. It can be observed that the use of SSFDA indeed shifts the distribution of accuracy across different subjects upward as a whole. Particularly, our method exhibits superior performance after applying SSFDA compared to the other two methods. Furthermore, the standard deviation of performance across various subjects is minimal for our method, demonstrating the robustness of our methodology when employing SSFDA. We detailed the differences by individuals in Figure 10(a) 10(b) in the appendix.
@@ A.1 Details of Data Collection
The experiments were carried out using the DataLITE wireless LE230 and DataLITE PIONEER, commercial sEMG acquisition systems from Biometrics Ltd.***https://www.biometricsltd.com/ The device’s sampling rate is 2000Hz, allowing for high-resolution data capture of the electrical activities in the muscles during the performance of gestures. Eight LE230 sEMG sensors were uniformly and equidistantly affixed to the surface of the participant’s right forearm.
@@ A.2 Details of Data Preprocessing
$\operatorname{RMS}\left(\mathbf{X}\right)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}x_{i}^{2}},$ (11)
@@ A.4 Conv-based Leaky Integrate-and-Fire Neurons
To transform the input into a spike train and learn the spike representation of sEMG samples, we use a smaller quantity of Conv-based Leaky Integrate-and-Fire (LIF) neurons. As shown in Figure 8(a), Conv-based LIF neurons can process spatial and temporal features concurrently. Since sEMG signals consist of pulses from multiple motor neurons at the skin’s surface combined with noise, using LIF neurons is biologically plausible for decoding these signals into spikes. Conv-LIF neurons also constrain the decoding range within local time, mitigating the impact of noise on global decoding. This method aligns with the biological characteristics of sEMG signals and effectively manages noise interference.
Initially, convolution is used to extract spatial features $S^{t,n}$ from the present spatial input $X^{t}$. These extracted features are then combined with the temporal features $T^{t-1,n}$ of the previous moment to form the current membrane potential $M^{t}$. Following this, the Fire and Leak module, based on the current membrane potential, generates the temporal features $T^{t,n}$ for the subsequent moment and the spatial features $S^{t,n+1}$for the next layer of the network. A significant event, known as a spike, occurs if the membrane potential $M^{t}$ exceeds a threshold value, denoted as $V_{th}$. In such an event, $M^{t}$ is reset to a value $V_{\text{reset}}$, and $T^{t,n}$, the post-synaptic potential, is calculated as the product of $S^{t,n+1}$ and $V_{\text{reset}}$. On the contrary, if the membrane potential $M^{t}$ is lesser than the threshold $V_{th}$, a spike does not occur. In this case, $T^{t,n}$ is determined by $M^{t}$ and is computed as the product of the time constant $e^{-dt/\tau}$ and $M^{t}$.
@@ A.6 Details of LIF-based Classifier
$\begin{cases}M^{t}=T^{t-1}+X^{t},\\ S^{t}=Hea\left(M^{t}-u_{\text{th}}\right),\\ T^{t}=V_{\text{reset}}S^{t}+\left(e^{-\frac{dt}{\tau}}M^{t}\right)\odot\left(1-S^{t}\right).\end{cases}$ (13)
In these equations, $M^{t}$ represents the membrane potential at the $t$-th time step. $X^{t}$ represents the spatial feature and $T^{t-1}$ represents the temporal feature. $\odot$ stands for element-wise multiplication. When $M^{t}$ reaches or exceeds the threshold $u_{\text{th}}$, it is reset to $V_{\text{reset}}$, and a spike ($S^{t}$ = 1) is emitted. Otherwise, $M^{t}$ evolves according to the given dynamics, and no spike is emitted ($S[n]=0$). $e^{-dt/\tau}<1$ stands for the decay factor.